OpenStudy (mrs.ambrose614):

@jigglypuff314 CAN YOU HELP ME WITH THIS PLEASE. Question 1 (Multiple Choice Worth 1 points) (03.02 LC) If h(x) = −5x − 10, find h(−2). −20 1.6 −1 0

OpenStudy (3mar):

Could I help you?

OpenStudy (calculusxy):

Substitue x for -2 in the function h(x) = -5x - 10 h(-2) = -5(-2) - 10 <- evaluate this and find the answer

OpenStudy (mrs.ambrose614):

Any one can help me can you help me out?????

OpenStudy (3mar):

Of course. Pleasure is mine! h(-2) means the value of the function h(x) at x=-2, so what you need to ir substitute with x=-2 in h(x) and find h(-2) In other words, replace every x in h(x)=-5x-10 with x=-2 this leads to: h(-2)=-5(-2)-10=?

OpenStudy (3mar):

Thank you for the medal @IrishBoy123! I do appreciate it.